A chemical supply company ships a certain solvent in 10-gallon drums. Let X represent the number of drums ordered by a randomly chosen customer. Assume X has the following probability mass function:
x 1 2 3 4 5
p(x) 0.33 0.2 0.2 0.17 0.1
Find the standard deviation of the number of drums ordered.

Answer by Boreal(15235)

The mean is x*p(x)=
0.33+0.4+0.6+0.68+0.5=2.51
The variance is sum (x-expected value or mean^2)*p(x)
(-1.51)^2*0.33+(2-2.51)^2*0.2+(3-2.51)^2*0.2+(4-2.51)^2*0.17+(5-2.51)^2*0.1=
0.752+0.052+0.048+0.377+0.620=1.849
sd is sqrt(1.849)=1.36 ANSWER